OpenStudy (anonymous):

Medal&Fan At the end of the winter, all skis at a sporting goods store were on sale for 3/5 off. What was the sale price of a snowboard that originally cost $360? Select one: a. $120 b. $240 c. $216 d. $144

OpenStudy (johnweldon1993):

Well if it was originally 360...and now on sale for 3/5 (or 60%) its original price... that means it NOW cost 2/5 of what it originally cost...so \[\large 360 \times \frac{2}{5} = \frac{720}{5} = 144\]

OpenStudy (anonymous):

@johnweldon1993 wow. thanks really! Its wonderful you actually show me how to do it.

OpenStudy (johnweldon1993):

Of course! Anytime :)